The viral load of COVID-19 in a sample of blood follows a normal distribution. The average number of viral particles is 500,​​​​​​​please show all work
Answer #1

Let X denotes the number of viral particles in a randomly selected person.

Here,

X ~ Normal(500, 302)

X - 500 P(X > 425) = 1 - P(X< 425) = 1-PC 30 425 – 500 30 = P(X > 425) = 1 - P(Z < 425 – 500 30 = P(X > 425) = 1- P(Z < -2.50